# MoORE: SVD-based Model MoE-ization for Conflict- and Oblivion-Resistant Multi-Task Adaptation

### v3

@@ Abstract
Adapting large-scale foundation models in multi-task scenarios often suffers from task conflict and oblivion. To mitigate such issues, we propose a novel “model MoE-ization” strategy that leads to a conflict- and oblivion-resistant multi-task adaptation method. Given a weight matrix of a pre-trained model, our method applies SVD to it and introduces a learnable router to adjust ithyds singular values based on tasks and samples. Accordingly, the weight matrix becomes a Mixture of Orthogonal Rank-one Experts (MoORE), in which each expert corresponds to the outer product of a left singular vector and the corresponding right one. We can improve the model capacity by imposing a learnable orthogonal transform on the right singular vectors. Unlike low-rank adaptation (LoRA) and its MoE-driven variants, MoORE guarantees the experts’ orthogonality and maintains the column space of the original weight matrix. These two properties make the adapted model resistant to the conflicts among the new tasks and the oblivion of its original tasks, respectively. Experiments on various datasets demonstrate that MoORE outperforms existing multi-task adaptation methods consistently, showing its superiority in terms of conflict- and oblivion-resistance. The code of the experiments is available at https://github.com/DaShenZi721/MoORE.
@@ 1 Introduction
Parameter-efficient adaptation/fine-tuning [64] plays a central role in the practical deployment of large-scale pre-training models, especially in multi-task learning (MTL) scenarios [73, 57, 11, 58, 40]. Take large language models (LLMs) [75] in natural language processing (NLP) tasks as an example. To increase intrinsic knowledge and maintain generalization power, a pre-trained LLM often needs to learn multiple downstream tasks in different domains simultaneously or sequentially.
To achieve multi-task adaptation, parameter-efficient fine-tuning (PEFT) methods like low-rank adaptation (LoRA) [24] and its variants [62, 2, 56, 60, 30] have been proposed. However, the practical applications of these methods are often limited because they suffer from task conflict and oblivion: $i$) The model adapted for one task often leads to the performance degradation in other tasks (also known as negative transfer [63, 72] or destructive interference [11]). $ii)$ Learning new tasks sometimes results in catastrophic forgetting [58] — the model performance deteriorates severely in previously learned tasks.
(a) An illustration of our SVD-based model MoE-ization strategy and MoORE architecture
(b) Resistance to task conflict
(c) Resistance to task oblivion
(d) Inference efficiency
Figure 1: (a) An illustration of our model MoE-ization strategy and the corresponding MoORE architecture. (b) The comparison for various multi-task adaptation methods on fine-tuning LLaMA-3.1 8B [20] on the CSR-MTL constructed by nine tasks [9, 8, 43, 5, 51, 70, 50, 55]. MoORE consistently works better than the baselines when the number of tasks is larger than one. (c) Before adaptation, LLaMA-3.1 8B achieves encouraging overall performance (i.e., the gray dashed line) in seven tasks [23, 22, 76, 54, 48, 7, 4, 10]. MoORE mitigates performance degradation and outperforms other baselines when the number of tasks exceeds three. (d) MoORE’s runtime is comparable to that of its competitors. Compared to the original LLaMA-3.1 8B (i.e., the gray dashed line), MoORE increases the inference time moderately.
@@
In principle, we call the above strategy “Model MoE-ization” because it converts the original neural network layer to an MoE architecture. In this study, we propose a novel model MoE-ization strategy, with the help of the singular value decomposition (SVD), leading to a conflict- and oblivion-resistant multi-task adaptation method. As illustrated in Figure 1(a), given a weight matrix of a pre-trained model, our method applies SVD [1] to it and introduces a learnable router to adjust its singular values based on tasks and samples. Accordingly, the weight matrix becomes a Mixture of Orthogonal Rank-one Experts (MoORE), in which each expert is constructed by the outer product of a left singular vector and the corresponding right one. To further improve the capacity of MoORE, we impose a learnable orthogonal adapter on the right singular vectors, which is implemented by a Householder reflection adaptation module [68].
Unlike existing methods, which learn some additional low-rank experts without any constraints on their relations, MoORE extracts many rank-one experts intrinsically from the SVD of the pre-trained weight matrix. Such a design guarantees the orthogonality of the experts, avoiding their information redundancy and undesired interferences across them. Moreover, similar to existing orthogonal fine-tuning strategies [46, 68], MoORE maintains the column space of the original weight matrix and thus makes the adapted model resistant to forgetting its original tasks. Thanks to the above two properties, MoORE consistently outperforms state-of-the-art methods in various multi-task adaptation scenarios. The representative results in Figure 1 show the superiority of MoORE in mitigating task conflict and oblivion and its competitive inference efficiency.
@@ 3.1 SVD-based Model MoE-ization
$&\bm{y}=\bm{WHx}^{(k)}+\sideset{}{{}_{d=1}^{D}}{\sum}\underbrace{g_{d}(\bm{x}^{(k)})}_{\text{router}}\underbrace{(\bm{u}_{d}\bm{v}_{d}^{\top}\bm{H})}_{\text{expert}}\bm{x}^{(k)}=\bm{U}\text{diag}(g(\bm{x}^{(k)})+\bm{\sigma})\bm{V}^{\top}\bm{H}\bm{x}^{(k)},\\ &\text{where}~{}~{}g(\bm{x}^{(k)})=\bm{P}^{\top}\bm{t}_{k}+\bm{Q}^{\top}\bm{\Gamma}\bm{x}^{(k)}\in\mathbb{R}^{D}.\end{aligned}$ (6)
@@
As shown in (7), the overall complexity of MoORE is $\mathcal{O}(D(D_{\text{out}}+D+D_{t}+D_{s}+L))$. In practice, we set $L,D_{t},D_{s}\ll\min\{D,D_{\text{out}}\}$ to reduce the complexity. Moreover, we can merge the orthogonal adapter into each expert in the inference phase, i.e., obtaining $\bm{V}^{\prime}=\bm{V}^{\top}\bm{H}$, and the complexity further reduces to $\mathcal{O}(D(D_{\text{out}}+D+D_{t}+D_{s}))$. Figure 1(d) shows that the inference efficiency of MoORE is comparable to its competitors.
@@
MTL-LoRA [66] $\text{Softmax}(\bm{\phi}_{k})$$\{\bm{B}_{m}\bm{\Lambda}_{k}\bm{A}\}_{m=1,k=1}^{M,K}$ $MK$ $r$ No No
@@
MoORE (Ours) $\bm{P}^{\top}\bm{t}_{k}+\bm{Q}^{\top}\bm{\Gamma}\bm{x}$$\{\bm{u}_{d}\bm{v}_{d}^{\top}\bm{H}\}_{d=1}^{D}$ $D$ $1$ Yes Yes
@@ 3.2 Comparisons with Existing MoE-based Multi-Task Adaptation Methods
  • The design of experts. Most existing methods apply $M$ low-rank adapters as experts, i.e., $\{\bm{B}_{m}\bm{A}_{m}\}_{m=1}^{M}$, where $\bm{B}\in\mathbb{R}^{D_{\text{out}}\times r}$ and $\bm{A}\in\mathbb{R}^{r\times D}$ are two rank-$r$ matrices. To reduce the number of learnable parameters, some methods, e.g., MoSLD [74], HydraLoRA [56], and MTL-LoRA [66], reuse the same $\bm{A}$ or $\bm{B}$ for all experts. MTL-LoRA [66] further introduces a task-specific matrix $\bm{\Lambda}_{k}$ for each expert, where $k=1,...,K$ indicates the task index. As a result, it creates $MK$ low-rank experts and activates $M$ experts per task. Our MoORE contains $D$ orthogonal rank-one experts $\{\bm{u}_{d}\bm{v}_{d}^{\top}\bm{H}\}_{d=1}^{D}$, in which the orthogonal adapter $\bm{H}$ is shared by all experts. Unlike existing methods, MoORE applies many simple but structured experts. Such a design has several advantages:
@@
      Imposing orthogonality for mitigating task conflict: The experts of MoORE are orthogonal to each other because $\bm{u}_{d}^{\top}\bm{u}_{d^{\prime}}=0$ for all $d\neq d^{\prime}$. The orthogonality ensures that the experts have different functionalities and domain knowledge, without redundant information. In particular, by activating different experts for different tasks, MoORE suppresses the interferences across the tasks in the training phase, thus mitigating the task conflict issue. Among existing methods, OMoE [18] is the only one imposing orthogonality on experts. However, it applies Gram-Schmidt orthogonalization algorithm [6] to the concatenation of $M$ experts’ output, i.e., $\text{GS}([\bm{B}_{1}\bm{A}_{m}\bm{x},\cdots,\bm{B}_{M}\bm{A}_{m}\bm{x}])$. As a result, imposing orthogonality requires additional $\mathcal{O}(D_{\text{out}}M^{2})$ operations per sample, which is less efficient than ours.
@@
LoRAMoE [14] $MD$ $M(D_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}D+M(D+D_{\text{out}}r+Dr))$
MixLoRA [29] $MD$ $M(D_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}D+MD+(D_{\text{out}}r+Dr))$
MoSLD [74] $MD$ $(D_{\text{out}}+MD)r$ $\mathcal{O}(D_{\text{out}}D+D_{\text{out}}r+M(Dr+D))$
HydraLoRA [56] $MD$ $(MD_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}D+Dr+M(D_{\text{out}}r+D))$
MTL-LoRA [66] $MK$ $(MD_{\text{out}}+Kr+D)r$ $\mathcal{O}(D_{\text{out}}D+Dr+MD_{\text{out}}r+r^{2})$
OMoE [18] $MD$ $M(D_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}(D+M^{2})+M(D+D_{\text{out}}r+Dr))$
MoORE (Ours) $D_{t}K+(D_{t}+2D_{s})D$ $LD$ $\mathcal{O}((D_{\text{out}}+D+D_{s}+D_{t}+L)D)$
@@ 4 Experiments
To demonstrate the effectiveness of MoORE in multi-task adaptation, we apply three MTL datasets and conduct comprehensive experiments on them. Representative results are shown in Figure 1 and the following content. More experimental details, e.g., the basic information of datasets, hyperparameter settings, ablation studies, routing weight analysis, and numerical results associated with figures, are shown in the Appendix.
@@ 4.1 Implementation Details
Base model and baselines. In the following experiments, we utilize LLaMA-3.1 8B [^2] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct [20] as the base model and adapt it by various multi-task adaptation methods. In particular, we compare MoORE with LoRA [24] and the methods incorporating low-rank adapters as MoEs, including LoRAMoE [14], MoSLD [74], MTL-LoRA [66], HydraLoRA [56], and MixLoRA [29]. We implement the MoE architectures of the baselines mainly based on their default settings. For a fair comparison, we modify some baselines’ hyperparameters to make the number of learnable parameters comparable for each method. For MoORE, we MoE-ize all linear layers of LLaMA-3.1 8B, including the “$QKVO$” modules of attention layers and the weight matrices of FFN layers.
@@
One dataset for evaluating oblivion-resistance. In addition, we construct one more dataset, called OR-MTL, for evaluating the oblivion-resistance of different methods. The dataset includes seven tasks, including MMLU [23, 22], IFEval [76], BIG-Bench Hard (BBH) [54], GPQA [48], HumanEval [7], MBPP [4], and GSM-8K [10]. The base model, LLaMA-3.1 8B, can achieve encouraging performance in these tasks. After adapting it on CSR-MTL, we record the performance of the adapted models in OR-MTL and assess the ability of different adaptation methods to mitigate task oblivion.

### v4

@@ Abstract
Adapting large-scale foundation models in multi-task scenarios often suffers from task conflict and oblivion. To mitigate such issues, we propose a novel “model MoE-ization” strategy that leads to a conflict- and oblivion-resistant multi-task adaptation method. Given a weight matrix of a pre-trained model, our method applies SVD to it and introduces a learnable router to adjust its singular values based on tasks and samples. Accordingly, the weight matrix becomes a Mixture of Orthogonal Rank-one Experts (MoORE), in which each expert corresponds to the outer product of a left singular vector and the corresponding right one. We can improve the model capacity by imposing a learnable orthogonal transform on the right singular vectors. Unlike low-rank adaptation (LoRA) and its MoE-driven variants, MoORE guarantees the experts’ orthogonality and maintains the column space of the original weight matrix. These two properties make the adapted model resistant to the conflicts among the new tasks and the oblivion of its original tasks, respectively. Experiments on various datasets demonstrate that MoORE outperforms existing multi-task adaptation methods consistently, showing its superiority in terms of conflict- and oblivion-resistance. The code is available at https://github.com/DaShenZi721/MoORE.
@@ 1 Introduction
Parameter-efficient adaptation/fine-tuning [71] plays a central role in the practical deployment of large-scale pre-training models, especially in multi-task learning (MTL) scenarios [80, 61, 11, 62, 42]. Take large language models (LLMs) [82] in natural language processing (NLP) tasks as an example. To increase intrinsic knowledge and maintain generalization power, a pre-trained LLM often needs to learn multiple downstream tasks in different domains simultaneously or sequentially. To achieve multi-task adaptation, parameter-efficient fine-tuning (PEFT) methods like low-rank adaptation (LoRA) [25] and its variants [68, 2, 60, 65, 31] have been proposed. However, the practical applications of these methods are often limited because they often suffer from task conflict and oblivion: $i$) The model adapted for one task often leads to the performance degradation in other tasks (also known as negative transfer [70, 79] or destructive interference [11]). ii)ii) Learning new tasks may result in catastrophic forgetting [62] — the model performance deteriorates severely in previously learned tasks.
(a) (a) An illustration of our model MoE-ization strategy and the corresponding MoORE architecture. (b) The comparison for various multi-task adaptation methods on fine-tuning LLaMA-3.1 8B [20] on the CSR-MTL constructed by nine tasks [9, 8, 45, 5, 54, 77, 53, 59]. MoORE consistently works better than the baselines when the number of tasks is larger than one. (c) Before adaptation, LLaMA-3.1 8B achieves encouraging overall performance (i.e., the gray dashed line) in seven tasks [24, 23, 83, 58, 51, 7, 4, 10]. MoORE mitigates the performance degradation and outperforms the most competitive baseline LoRAMoE [14]. (d) MoORE’s runtime is comparable to that of its competitors. Compared to the original LLaMA-3.1 8B (i.e., the gray dashed line), MoORE increases the inference time moderately.
@@
In principle, we call the above strategy “Model MoE-ization” because it converts the original linear layer to an MoE architecture. In this study, we propose a novel model MoE-ization strategy, with the help of the singular value decomposition (SVD), leading to a conflict- and oblivion-resistant multi-task adaptation method. As illustrated in Figure LABEL:fig:scheme, given a weight matrix of a pre-trained model, our method applies SVD [1] to it and introduces a learnable router to adjust its singular values based on tasks and samples. Accordingly, the weight matrix becomes a Mixture of Orthogonal Rank-one Experts (MoORE), where each expert is constructed by the outer product of a left singular vector and the corresponding right one. To further improve the capacity of MoORE, we impose a learnable orthogonal adapter, which is implemented by a Householder reflection adaptation module [75].
Unlike existing methods, which learn some additional low-rank experts without any constraints on their relations, MoORE extracts many rank-one experts intrinsically from the SVD of the original weight matrix. Such a design imposes the orthogonality of the experts, avoiding their information redundancy and undesired interferences. Moreover, similar to existing orthogonal fine-tuning strategies [49, 75], MoORE maintains the column space of the original weight matrix and thus inherit the pre-training ability better. Thanks to the above two properties, MoORE consistently outperforms SOTA methods in various multi-task adaptation scenarios. The results in Figure 4(a) show the superiority of MoORE in mitigating task conflict and oblivion and its competitive inference efficiency.
@@ 3.1 SVD-based Model MoE-ization
$\begin{aligned} &\bm{y}=\bm{WHx}^{(k)}+\sideset{}{{}_{d=1}^{D}}{\sum}\underbrace{g_{d}(\bm{x}^{(k)})}_{\text{router}}\underbrace{(\bm{u}_{d}\bm{v}_{d}^{\top}\bm{H})}_{\text{expert}}\bm{x}^{(k)}=\bm{U}\text{diag}(g(\bm{x}^{(k)})+\bm{\sigma})\bm{V}^{\top}\bm{H}\bm{x}^{(k)},\\ &\text{where}~~g(\bm{x}^{(k)})=\bm{P}^{\top}\bm{t}_{k}+\bm{Q}^{\top}\bm{\Gamma}\bm{x}^{(k)}\in\mathbb{R}^{D}.\end{aligned}$ (6)
@@
As shown in (7), the overall complexity of MoORE is $\mathcal{O}(D(D_{\text{out}}+D+D_{t}+D_{s}+L))$. In practice, we set $L,D_{t},D_{s}\ll\min\{D,D_{\text{out}}\}$ to reduce the complexity. Moreover, we can merge the orthogonal adapter into each expert in the inference phase, i.e., obtaining $\bm{V}^{\prime}=\bm{V}^{\top}\bm{H}$, and the complexity further reduces to $\mathcal{O}(D(D_{\text{out}}+D+D_{t}+D_{s}))$. Figure LABEL:fig:speed shows that the inference efficiency of MoORE is comparable to its competitors.
@@
MTL-LoRA [73] $\text{Softmax}(\bm{\phi}_{k})$ $\{\bm{B}_{m}\bm{\Lambda}_{k}\bm{A}\}_{m=1,k=1}^{M,K}$ MKMK $r$ No No
@@
MoORE (Ours) $\bm{P}^{\top}\bm{t}_{k}+\bm{Q}^{\top}\bm{\Gamma}\bm{x}$ $\{\bm{u}_{d}\bm{v}_{d}^{\top}\bm{H}\}_{d=1}^{D}$ $D$ 11 Yes Yes
@@ 3.2 Comparisons with Existing MoE-based Multi-Task Adaptation Methods
  • The design of experts. Most existing methods apply $M$ low-rank adapters as experts, i.e., $\{\bm{B}_{m}\bm{A}_{m}\}_{m=1}^{M}$, where $\bm{B}\in\mathbb{R}^{D_{\text{out}}\times r}$ and $\bm{A}\in\mathbb{R}^{r\times D}$ are two rank-$r$ matrices. To reduce the number of learnable parameters, some methods, e.g., MoSLD [81], HydraLoRA [60], and MTL-LoRA [73], reuse the same $\bm{A}$ or $\bm{B}$ for all experts. MTL-LoRA [73] further introduces a task-specific matrix $\bm{\Lambda}_{k}$ for each expert, where k=1,,Kk=1,...,K indicates the task index. As a result, it creates $MK$ low-rank experts and activates $M$ experts per task. Our MoORE contains $D$ orthogonal rank-one experts $\{\bm{u}_{d}\bm{v}_{d}^{\top}\bm{H}\}_{d=1}^{D}$, in which the orthogonal adapter $\bm{H}$ is shared by all experts. Unlike existing methods, MoORE applies many simple but structured experts. Such a design has several advantages:
@@
      Imposing orthogonality for mitigating task conflict: The experts of MoORE are orthogonal to each other because $\bm{u}_{d}^{\top}\bm{u}_{d^{\prime}}=0$ for all $d\neq d^{\prime}$. The orthogonality ensures that the experts have different functionalities and domain knowledge, without redundant information. In particular, by activating different experts for different tasks, MoORE suppresses the interferences across the tasks in the training phase, thus mitigating the task conflict issue. It is worth noting that O-LoRA [67] also introduces orthogonality constraints within its LoRA adapters. However, there are two notable differences between O-LoRA and MoORE: $i)$ O-LoRA uses a regularization term to pursue orthogonality, which can not strictly make its LoRA adapters orthogonal to each other. ii)ii) O-LoRA does not adopt a MoE architecture. To deal with different tasks, O-LoRA adds new LoRA adapters for each sequentially incoming task, rather than designing and learning a flexible routing mechanism for given and fixed experts. As a result, the model complexity of O-LoRA is linear to the number of tasks. Among existing MoE-based methods, OMoE [18] is the only one imposing orthogonality on experts. However, it applies Gram-Schmidt orthogonalization algorithm [6] to the concatenation of $M$ experts’ output, i.e., $\text{GS}([\bm{B}_{1}\bm{A}_{m}\bm{x},\cdots,\bm{B}_{M}\bm{A}_{m}\bm{x}])$. As a result, imposing orthogonality requires additional $\mathcal{O}(D_{\text{out}}M^{2})$ operations per sample, which is less efficient than ours.
@@
LoRAMoE [14] MDMD $M(D_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}D+M(D+D_{\text{out}}r+Dr))$
MixLoRA [30] MDMD $M(D_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}D+MD+(D_{\text{out}}r+Dr))$
MoSLD [81] MDMD $(D_{\text{out}}+MD)r$ $\mathcal{O}(D_{\text{out}}D+D_{\text{out}}r+M(Dr+D))$
HydraLoRA [60] MDMD $(MD_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}D+Dr+M(D_{\text{out}}r+D))$
MTL-LoRA [73] MKMK $(MD_{\text{out}}+Kr+D)r$ $\mathcal{O}(D_{\text{out}}D+Dr+MD_{\text{out}}r+r^{2})$
OMoE [18] MDMD $M(D_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}(D+M^{2})+M(D+D_{\text{out}}r+Dr))$
MoORE (Ours) $D_{t}K+(D_{t}+2D_{s})D$ LDLD $\mathcal{O}((D_{\text{out}}+D+D_{s}+D_{t}+L)D)$
3.3 Comparisons with Existing SVD-based Fine-Tuning Methods
Table 3: Comparisons for various SVD-based fine-tuning methods on their implementations.
Method Model in Training Learnable Parameters Model in Inference
SVF [56] $\bm{U}\text{diag}(\sigma)\bm{V}^{\top}$ Nonparametric $\sigma$ Original architecture: $\bm{W}_{\text{new}}$
SVDiff [21] $\bm{U}\text{diag}(\text{ReLU}(\sigma+\delta))\bm{V}^{\top}$ Nonparametric $\delta$ Original architecture: $\bm{W}_{\text{new}}$
SVFT [33] $\bm{U}(\text{diag}(\sigma)+\bm{M})\bm{V}^{\top}$ Nonparametric $\bm{M}$ Original architecture: $\bm{W}_{\text{new}}$
KaSA [64] $\bm{W}_{lr}+\Delta\bm{U}\text{diag}(\Delta\sigma)\Delta\bm{V}^{\top}$ Nonparametric $\Delta\bm{U},\Delta\sigma,\Delta V^{\top}$ Original architecture: $\bm{W}_{\text{new}}$
SVFCL [69] $\bm{U}\text{diag}(\sigma+\sum_{i=0}^{t}\sigma_{i})\bm{V}^{\top}$ Nonparametric $\sigma_{i}$’s Original architecture: $\bm{W}_{\text{new}}$
MoORE (Ours) $\bm{WH}$ $\sum_{d=1}^{D}g_{d}(\bm{x}^{(k)})(\bm{u}_{d}\bm{v}_{d}^{\top}\bm{H})$ Parametric $g(\cdot)$ and Householder reflections $\bm{H}$ An MoE, where $\bm{WH}$ and
$\bm{v}_{d}^{\top}\bm{H}$ are pre-computed
Table 4: Comparisons for various SVD-based fine-tuning methods on their applications.
Method Multi-Task Learning Consider Task Conflict Consider Task Oblivion
SVF [56] No No No
SVDiff [21] No No No
SVFT [33] No No No
KaSA [64] No No No
SVFCL [69] Yes (incremental learning) No Partially discussed
MoORE (Ours) Yes Yes, by orthogonal experts Yes, by maintaining $\text{Range}(\bm{W})$
Our method employs SVD to perform model MoE-ization, constructing a complete MoE architecture. This fundamentally differs from existing SVD-based fine-tuning methods [48, 56, 21, 33, 64, 69]. To more clearly illustrate the differences between MoORE and these methods, we select five recent ones for comparison, including SVF [56], SVDiff [21], SVFT [33], KaSA [64], and SVFCL [69]. Tables 3 and 4 compare different methods in terms of implementation and application, respectively.
  • In the aspect of architecture, MoORE is the only method that learns a parametric router to adjust singular values, leading to an MoE architecture for inference. In MoORE, each expert is constructed by the outer product of singular vectors and the learnable Householder reflections.
  • In the aspect of application, only MoORE and SVFCL consider multi-task adaptation. MoORE is applicable for both simultaneous and sequential multi-task adaptation, while SVFCL only considers incremental learning. Moreover, MoORE makes the first attempt to connect task conflict and oblivion to the orthogonality of experts and the column space of the parameter matrix.
@@ 4 Experiments
To demonstrate the effectiveness of MoORE in multi-task adaptation, we apply three MTL datasets and conduct comprehensive experiments on them. Representative results are shown in Figure 4(a) and the following content. More experimental details, e.g., the basic information of datasets, hyperparameter settings, ablation studies, routing weight analysis, and numerical results associated with figures, are shown in the Appendix.
@@ 4.1 Implementation Details
Base model and baselines. In the following experiments, we utilize LLaMA-3.1 8B [^2] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct [20] as the base model and adapt it by various multi-task adaptation methods. In particular, we compare MoORE with LoRA [25] and the methods incorporating low-rank adapters as MoEs, including LoRAMoE [14], MoSLD [81], MTL-LoRA [73], HydraLoRA [60], and MixLoRA [30]. We implement the MoE architectures of the baselines mainly based on their default settings. For a fair comparison, we modify some baselines’ hyperparameters to make the number of learnable parameters comparable for each method. For MoORE, we MoE-ize all linear layers of LLaMA-3.1 8B, including the “QKVOQKVO” modules of attention layers and the weight matrices of FFNs.
@@
One dataset for evaluating oblivion-resistance. In addition, we construct one more dataset, called OR-MTL, for evaluating the oblivion-resistance of different methods. The dataset includes seven tasks, including MMLU [24, 23], IFEval [83], BIG-Bench Hard (BBH) [58], GPQA [51], HumanEval [7], MBPP [4], and GSM-8K [10]. The base model, LLaMA-3.1 8B, achieves encouraging performance in these tasks. After adapting it on CSR-MTL, we record the performance of the adapted models in OR-MTL and assess the ability of different methods to mitigate task oblivion.